• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nanda

    Log entry time 21:02:29 on June26,2004

    Entry number 127012

    keyword=compton controls

    Compton control problems. Laser and photon table stopped responding to EPICS. Rebooting IOC fixed the Laser problems and cavity locked. But the X-axis of photon table remained stuck.

    Made access to the Hall to reset motion controller. The x-axis still did not reset. Moved the table manually to align central crystal to the compton photons (eyeball precision!). This might be ok for now to go on with data taking until we find a solution to the controller problem.